X Company currently buys 6,000 units of a part each year from a supplier for $8.30 per part, but it is considering making the part instead. In order to make the part, X Company will have to buy equipment that will cost $150,000. The equipment will last for 6 years, at which time it will have zero disposal value. X Company estimates that it will cost $20,250 a year to make all 6,000 units.

What is the approximate rate of return if X Company makes the part instead of buying it from the supplier? [Note: 0.03 means 3%, etc.]
Answer #1

Annual Depreciation on equipment = $150000 / 6 = $25,000

Total relevant cost of making the product = 25000 + 20250 = $45,250

Return on making the product insted of buying = 6000*8.30 - 45250 = $4550

Rate of return = Return on making the product insted of buying / Investment

= 4550 / 1500000

= 3%